OpenStudy (anonymous):

Solve with this point slope formula y-y1=m(x-x1) (-8,7);(8,15).Can someone Please Show Steps and Instructions and explaining on how to do it Please??

OpenStudy (jlastino):

You first get the slope using the formula (y2-y1)/(x2-x1) where x1=-8 x2=8 y1=7 y2=15 This would be the "m" in your equation Then sub the values into your equation

OpenStudy (anonymous):

So First Use The Point Slope Formula. Then??

OpenStudy (jlastino):

Get the slope first. Then use that slope in the point slope formula

OpenStudy (anonymous):

And then?

OpenStudy (anonymous):

What Would the Answer Be??

OpenStudy (jlastino):

y=1/2x +11

OpenStudy (jlastino):

I mean y= (1/2)x +11

OpenStudy (anonymous):

so which formula would i use the one i wrote in the question?

OpenStudy (jlastino):

yeah it would result in what i just wrote

OpenStudy (anonymous):

oh thanx alot

OpenStudy (jlastino):

you can also use the two point formula,